# Nearly a Decade of Groundbreaking Speckle Interferometry at the International Gemini Observatory

### v1

@@ Abstract
Since its inception, speckle interferometry has revolutionized high-resolution astronomical imaging, overcoming atmospheric challenges to achieve the diffraction limits of telescopes. Almost a decade ago, in 2018, a pair of speckle cameras – ‘Alopeke and Zorro – were installed at the twin 8.1-meter Gemini North and South telescopes, two of the largest apertures in the world, in Hawai’i and Chile. Equipped with dual blue and red channels, ’Alopeke and Zorro deliver high-resolution imaging across optical bandpasses from 350 to 1000 nm, which has led to crucial discoveries in both stellar multiplicity and exoplanetary science. Furthermore, the broad and nonrestrictive access to these instruments, given by each Gemini Observatory partner and via the US NOIRLab open skies policy, has allowed our community to expand the applications of the instruments, supporting a wide range of scientific investigations from Solar System bodies, to morphological studies of stellar remnants, to evolved stars, to transient phenomena. This paper reviews the instrument technology and observational capabilities, and highlights key scientific contributions and discoveries of ‘Alopeke and Zorro, emphasizing the enduring importance of speckle interferometry in advancing modern observational astronomy and expanding the frontiers of astronomical research.
Binary Stars (154), Direct Imaging (387), Exoplanet Detection (489), Exoplanets (498), Planet Hosting Stars (1242), Speckle Interferometry (1552), Astronomical Techniques (1684), High Angular Resolution (2167), High Contrast Techniques (2369)
journal: PASP
@@ 1 Introduction
Speckle interferometry, as a technique for high-resolution optical imaging, began in 1970 with the work of Labeyrie (1970), using ideas put forward mainly by Fried (1966). Labeyrie (1970) showed that taking short exposures removed the effects of seeing-induced fluctuations that were causing distortions in the wavefront from a distant star. Removing these distortions allowed the diffraction limit of the telescope to be reached.
Speckle cameras initially used photographic plates, but continued with the various detectors of the day, such as photomultiplier tubes, video tubes, and reticons (e.g., Bonneau & Foy, 1980; McAlister et al., 1987; Weigelt & Baier, 1985; Horch et al., 1992; Balega et al., 1993). More recently, speckle image reconstructions using digital data and Fourier analysis have produced more detailed image reconstructions (e.g., Howell et al., 2011; Hope et al., 2022). While initial studies mainly focused on bright binary stars using 1- to 4-meter telescopes and repeated imaging to produce precise stellar orbits (e.g. McAlister et al., 1989), Fourier-based data reduction and image reconstructions (Horch et al., 2012, 2015) are enhanced using the bispectrum technique developed by Weigelt (1977) and Lohmann et al. (1983). This technique obtains phase information, and the usual Fourier 180 degree ghost ambiguity is overcome.
Significant advances in astronomical detectors, in particular the more quantum efficient charge-coupled devices (CCDs), provided the next leap forward in this field, allowing for photon intensification, digital outputs, and higher signal-to-noise (S/N) observations to be obtained (e.g., McAlister et al., 1989; Mason et al., 1997; Hartkopf et al., 2000; Horch et al., 2000). In recent years, the introduction of the electron-multiplying CCD (EMCCD) as a detector (Tokovinin & Cantarutti, 2008; Horch et al., 2011) has been a game-changer. With ultra-fast readout, essentially zero read noise, near-perfect quantum efficiency, optical flatness, and ease of use, EMCCDs have revolutionized the field of speckle imaging.
This renaissance in speckle imaging has led to the development of new dedicated speckle instruments (e.g., Maksimov et al., 2009; Tokovinin et al., 2010; Howell et al., 2021d; Clark et al., 2020; Pedichini et al., 2016) placed on some of the largest telescopes in the world. A summary of former and current speckle imagers in astronomy is presented in Scott et al. (2021). Speckle imaging is no longer limited to bright star astrometry; it has expanded into many areas of point source and non-point source imagery (e.g., Salinas et al., 2020; Scott et al., 2021; Shara et al., 2022). Fainter astronomical targets can now be observed (Howell et al., 2021a), the overall data quality and S/N ratio of the observations are greater, and the final reconstructed images have more fidelity (e.g., Howell & Furlan, 2022a).
@@
This paper presents a summary of the first eight years of astronomical imaging observations using the highest resolution, deepest contrast speckle instruments available, ‘Alopeke and Zorro, which are mounted on the twin 8.1-meter Gemini North and South telescopes in Hawai’i and Chile (Scott et al., 2021; Howell & Furlan, 2022a). We review the major advances and scientific results obtained using these instruments since their introduction in 2018.
@@ 2 Zorro and ‘Alopeke: visiting speckle instruments at Gemini
‘Alopeke and Zorro (the ‘$\rm{\overline{o}}$lelo Hawai’i and Spanish words for “fox”) are identical instruments that use iXon Ultra 888 EMCCD cameras to provide simultaneous speckle imaging in two band passes, yielding high-resolution reconstructed images of the observed sources. These imagers are used with one of two fields-of-view: speckle mode ($6.7\times 6.7$ arcseconds) or wide-field mode ($60\times 60$ arcseconds). The standard data reduction pipeline (Horch et al., 2001; Howell et al., 2011) provides robust $5\sigma$ magnitude contrast limits on stellar companion or circumstellar material detections (e.g., Howell et al., 2016). The left panel of Figure 1 shows ‘Alopeke with the covers removed, revealing the tightly packed innards that contain two filter wheels, the optical elements, and the two ANDOR EMCCD cameras extending from the box. The right panel of Figure 1 shows Zorro in its permanent mount location, attached underneath the Gemini South 8.1-meter primary mirror along the calibration unit port. A more detailed description of these two visiting Gemini instruments can be found in Scott et al. (2021).
Figure 1: Left: The ‘Alopeke main instrument components were visible during a routine maintenance inspection at the Gemini North base facility. The ANDOR EMCCD cameras are seen extending from the main enclosure. Right: Zorro mounted on the Gemini South telescope. The black box contains the instrument, while the white box contains the electronic power supplies and instrument computer.
‘Alopeke and Zorro proposal demand at Gemini varies semester by semester, but on average it is 5-10%, exceeding GSAOI demand and nearing the demand for other visiting instruments (e.g., IGRINS), or even facility instruments (e.g., GNIRS). Any user can request time on these instruments using the regular Call for Proposals, or Director’s Discretionary Time (DDT) when needed. Principal Investigators (PIs) from Gemini partner countries can also propose for ‘Alopeke and Zorro time using Fast Turnaround (FT) proposals (depending on the Gemini instrument schedule [^1] https://www.gemini.edu/observing/schedules-and-queue/).
@@ 3 Observations, Data Reduction, and Image Reconstruction
Speckle instrumentation is fairly simple, inexpensive, and small compared to other instruments mounted on large-aperture telescopes. ‘Alopeke and Zorro are roughly the size of a carry-on suitcase. The most difficult aspect of speckle observing is the quantity and short length of the exposures, which is managed easily by the fast readout capabilities of the EMCCDs. Speckle imaging of a target requires many thousands of short exposures (10 to 60 milliseconds in length) to be obtained and processed. This large number of images is required to build up sufficient S/N, especially at contrasts greater than four or five magnitudes ($\sim$10-2), at very close angular separations, and/or for fainter targets. Although this large number of exposures may seem daunting, at tens of milliseconds per image, typical speckle observations last only a few minutes per target (Hope et al., 2019; Howell et al., 2021d). Stars with $V$ magnitudes from 1 to $\sim 12$ require only $\sim$5 minutes of observation time, during which 3,000 to 5,000 thousand exposures are collected. However, targets as faint as $R=19$ can be observed using $\sim$ 50 minutes of on-source time. In a typical night, 40 to 50 sources are observed. Scott et al. (2021) and Howell & Furlan (2022b) provide additional details on the relation between source magnitude, filter, and sky conditions to total on-source time. In Section 4, we list the total on-source times and filters used for each observation.
@@
Nearly all currently used speckle image reconstruction software packages are based on Fourier speckle interferometric methods (e.g., Labeyrie, 1970; Lohmann et al., 1983; Horch et al., 2001, 2015). The results presented in this paper and all fully reduced data in the archives are based on our implementation of these methods as described in Howell et al. (2011) and Horch et al. (2012). Fourier speckle deconvolution techniques are computationally efficient but require the routine observation of stars taken from the HR or HD catalogs that are known or assumed to be single based on past spectroscopic or imaging observations. These are called PSF standards, and are assigned to each target by the imaging team prior to the scheduled observations. They are observed near in time to the targets and with similar sky locations. Each PSF standard observation requires an additional $\sim 3$ minutes of observing time. Modern image reconstruction methods that are based on blind deconvolution techniques and can reach deeper contrasts and provide more accurate astrophysical results are beginning to be used as well (e.g., Howell et al., 2024).
@@
Figure 2 presents a typical speckle imaging result for a point source (a star), and the discovery of a very close stellar companion that is 2.4 magnitudes fainter than the primary (Scott et al., 2021; Howell & Furlan, 2022a). Additional raw and fully-reduced archival data from Zorro and ‘Alopeke can be found in the Gemini Observatory (Hirst & Cardenes, 2017) [^2] https://archive.gemini.edu/searchform/ZORRO,
@@
Figure 2: A typical speckle imaging result for an 8th magnitude star. The red and blue curves show the $5\sigma$ contrasts achieved as a function of angular separation, from the diffraction limit out to 1.2 arcseconds. The star symbol shows the properties of the companion, and the inset shows the reconstructed image for the 832 nm observation; a scale bar is included for reference. Our speckle observations revealed that the target star is binary, with a very close (0.06 arcsecond separation) companion that is 2.4 magnitudes fainter than the primary.
@@ 4 A Panoply of Scientific Applications and Results
In this section, we discuss various scientific investigations that have been carried out using observations from ‘Alopeke and/or Zorro. Most of these observations were obtained as test cases or engineering studies (unless otherwise noted) and as such, many have not been published previously. These studies allowed researchers to assess the potential for using speckle imaging to accomplish their scientific goals, and have formed the basis for a number of stand-alone proposals targeting the detailed study of single objects, or larger sample collection to enable more global results.
Figure 3: The microlens source Kojima-1Lb observed with ‘Alopeke at Gemini North. This SDSS $i$ band image was obtained using 35 minutes of on-source time and detected both the lens and source separated by 0.058 arcsec with a magnitude difference of 3.7.
Figure 4: Left: The late M star WISE J133734.15-371100.2 as observed by Zorro. This 9th magnitude star was observed for 30 minutes at 832 nm. Note the close companion at a separation of 0.6 arcsec that is 3 magnitudes fainter. At a distance of 34 pc, the two stars are 20 au apart and have masses of 0.23 and 0.09 M respectively. Right: The planet hosting star TOI-697 was discovered to be triple using Gemini speckle observations. The primary star harbors a 2.6 R planet in a 8.6 day orbit and is itself orbited by a close binary pair 5 magnitudes fainter at a separation of 1.15 arcsec. This 832 nm observation used 5 minutes of Gemini time.
@@ 4.1 Stars
Stars have been the predominant targets of our speckle imaging campaigns. In this section, we explore the varied results from these observations.
@@ 4.1.1 Exoplanet-Hosting Stars
Several large observing programs aimed at surveying exoplanet host stars have been in place at Gemini for many years. These programs have targeted transiting exoplanet candidates from Kepler, K2, TESS, and other NASA missions (e.g., Lester et al., 2021; Howell et al., 2021b, c; Clark et al., 2022; Matson et al., 2025). Such surveys aim to identify close-in stellar companions that either induce false positive signals or contaminate the light curves of planets, leading to underestimated planetary radii (Ciardi et al., 2015; Furlan & Howell, 2017). Figure 2 is an example of an 8th magnitude TESS exoplanet host star observed at Gemini South with Zorro. Observations of exoplanet host stars continue for missions such as TESS, Roman, Ariel, and the Habitable Worlds Observatory. As the Roman space telescope prepares for launch, microlensing observations will reach a new high as numerous free-floating planets will be discovered via observations pointed toward the Galactic bulge. Speckle observations will play a large role in resolving the lens from the target star, as was done for the microlens event Kojima-1Lb (Figure 3).
@@ 4.1.2 Stellar Multiplicity
Surveys investigating the multiplicity of various types of stars have been carried out at Gemini Observatory. These include low-mass stars (e.g., Winters et al., 2021; Clark et al., 2022), halo stars (e.g., Hartman et al., 2022), and higher-order multiplicity in known binaries (e.g., Tokovinin, 2023). Figure 4 shows the faint L-dwarf companion to an M star, as well as a newly discovered, planet-hosting triple star system. High-resolution imaging of triple star systems has also been proposed as a robust test of modified gravity (Manchanda et al., 2023).

### v2

@@ Abstract
1
Since its inception, speckle interferometry has revolutionized high-resolution astronomical imaging, overcoming atmospheric challenges to achieve the diffraction limits of telescopes. Almost a decade ago, in 2018, a pair of speckle cameras – ‘Alopeke and Zorro – were installed at two of the largest apertures in the world, the twin 8.1-meter Gemini North and South telescopes in Hawai’i and Chile. Equipped with dual blue and red channels, ’Alopeke and Zorro deliver high-resolution imaging across the optical bandpass from 350 to 1000 nm, which has led to crucial discoveries in both stellar multiplicity and exoplanetary science. Furthermore, the broad and nonrestrictive access to these instruments, given by each Gemini Observatory partner and via the US NOIRLab open skies policy, has allowed the community to expand the applications of the instruments, supporting a wide range of scientific investigations from Solar System bodies, to morphological studies of stellar remnants and quasars, to evolved stars, to transient phenomena. This paper reviews the instrument technology and observational capabilities, and highlights key scientific contributions and discoveries of ‘Alopeke and Zorro, emphasizing the enduring importance of speckle interferometry in advancing modern observational astronomy and expanding the frontiers of astronomical research.
@@
2 Keywords: Binary Stars, Planet Hosting Stars, Speckle Interferometry, Astronomical Techniques, High Angular Resolution
@@ 3 Introduction
Speckle interferometry, as a technique for high-resolution optical imaging, began in 1970 with the work of Labeyrie (1970), using ideas for fast imaging to remove atmospheric effects. Labeyrie (1970) showed that taking short exposures removed the effects of seeing-induced fluctuations that were causing distortions in the wavefront from a distant star. Removing these distortions allowed the diffraction limit of the telescope to be reached.
Speckle cameras initially used photographic plates, but continued with the various detectors of the day, such as photomultiplier tubes, video tubes, and reticons (e.g., Bonneau and Foy, 1980; McAlister et al., 1987; Weigelt and Baier, 1985; Horch et al., 1992; Balega et al., 1993). Initial studies mainly focused on bright binary stars using 1- to 4-meter telescopes and repeated imaging to produce precise stellar orbits (e.g. McAlister et al., 1989). Significant advances in astronomical detectors, in particular the more quantum efficient charge-coupled devices (CCDs), provided the next leap forward in this field, allowing for photon intensification, digital outputs, and higher signal-to-noise (S/N) observations to be obtained (e.g., McAlister et al., 1989; Mason et al., 1997; Hartkopf et al., 2000; Horch et al., 2000). In recent years, the introduction of the electron-multiplying CCD (EMCCD) as a detector (Tokovinin and Cantarutti, 2008; Horch et al., 2011) has been a game-changer. With ultra-fast readout, essentially zero read noise, near-perfect quantum efficiency, optical flatness, and ease of use, EMCCDs have revolutionized the field of speckle imaging. Fourier-based data reduction and image reconstructions using autocorrelation and power spectra techniques (Horch et al., 2012, 2015) are enhanced using the bispectrum technique developed by Weigelt (1977) and Lohmann et al. (1983) allowing phase information to be determined that resolves the ambiguity present in autocorrelation analysis (e.g., Howell et al., 2011; Hope et al., 2022).
This renaissance in speckle imaging has led to the development of new dedicated speckle instruments (e.g., Maksimov et al., 2009; Tokovinin et al., 2010; Howell et al., 2021b; Clark et al., 2020; Pedichini et al., 2016) placed on some of the largest telescopes in the world. A summary of former and current speckle imagers in astronomy is presented in Scott et al. (2021). Speckle imaging is no longer limited to bright star astrometry; it has expanded into many areas of point source and non-point source imagery (e.g., Salinas et al., 2020; Scott et al., 2021; Shara et al., 2022). Fainter astronomical targets can now be observed (Howell et al., 2021a), the overall data quality and S/N ratio of the observations are greater, and the final reconstructed images have more fidelity (e.g., Howell and Furlan, 2022).
@@
This paper presents a summary of the first eight years of astronomical imaging observations using the highest resolution, deepest contrast speckle instruments available, ‘Alopeke and Zorro, which are mounted on the twin 8.1-meter Gemini North and South telescopes in Hawai’i and Chile (Scott et al., 2021; Howell and Furlan, 2022). We review the major advances and scientific areas covered with these instruments since their introduction in 2018.
@@ 4 Zorro and ‘Alopeke: visiting speckle instruments at Gemini
[…]
Figure 1: Left: The ‘Alopeke main instrument components were visible during a routine maintenance inspection at the Gemini North base facility. The ANDOR EMCCD cameras are seen extending from the main enclosure. Right: Zorro mounted on the Gemini South telescope. The black box with one ANDOR camera seen extending from it contains the instrument, while the larger white box to the left contains the electronic power supplies and instrument computer.
‘Alopeke and Zorro proposal demand at Gemini varies semester by semester, but on average it is 5-10%, exceeding the Gemini South Adaptive Optics Imager (GSAOI) demand and nearing the demand for other visiting instruments (e.g., IGRINS), or even facility instruments (e.g., GNIRS). Any user can request time on these instruments using the regular Call for Proposals, or Director’s Discretionary Time (DDT) when needed. Principal Investigators (PIs) from Gemini partner countries can also propose for ‘Alopeke and Zorro time using Fast Turnaround (FT) proposals  [^2] https://www.gemini.edu/observing/schedules-and-queue/.
@@ 5 Observations, Data Reduction, and Image Reconstruction
Speckle instrumentation is fairly simple, inexpensive, and small compared to other instruments mounted on large-aperture telescopes. ‘Alopeke and Zorro are roughly the size of a carry-on suitcase. The most difficult aspect of speckle observing is the quantity and short duration of the exposures, which is easily managed by the fast read-out capabilities of the EMCCDs. Using the Gemini mirror coating reflectivity, the optics and filter transmissions, and the QE of the EMCCDs, optimal speckle observations can be carried out between 400-940 nm where the total throughput is 40% or higher (Scott et al., 2021). As a mitigation for the lack of an ADC, typical speckle observations are performed at an airmass of 1.4-1.5 or lower (elevations above 45 degrees) and narrow band filters are used. Speckle imaging at Gemini is often done under (bright) moonlit sky conditions, for which the narrow band filters are preferred as well. The EMCCDs can also function as normal CCDs, allowing traditional, long exposure digital images to be obtained.
@@
Speckle imaging of a target requires many thousands of short exposures (10 to 60 milliseconds in length) to be obtained and processed. This large number of images is required to build up sufficient S/N, especially at contrasts greater than four or five magnitudes ($\sim$10-2), at very close angular separations, and/or for fainter targets. Although this large number of exposures may seem daunting, at tens of milliseconds per image, typical speckle observations last only a few minutes per target (Hope et al., 2019; Howell and Furlan, 2022). Stars with $V$ magnitudes from 1 to $\sim 12$ require only $\sim$5 minutes of observation time, during which 3,000 to 5,000 thousand exposures are collected. However, targets as faint as $R=19$ can be observed by using $\sim$ 50 minutes of on-source time. In a typical night, 40 to 50 sources are observed. Scott et al. (2021) and Howell and Furlan (2022) provide full details on the relation between the magnitude of the source, the filter, the brightness of the target and the sky conditions to the total integration time at the source and the resulting S/N. To provide the reader with a gauge for the time required for a variety of observation types, we list in Section 6, the total on-source times and filters used for each observation.
Almost all currently used speckle image reconstruction software packages are based on Fourier speckle interferometric methods (e.g., Labeyrie, 1970; Lohmann et al., 1983; Horch et al., 2001, 2015). Our standard data reduction pipeline (Horch et al., 2012; Howell et al., 2011) provides robust $5\sigma$ magnitude contrast limits on stellar companion or circumstellar material detections (e.g., Howell et al., 2016). The results presented in this paper, and all fully reduced data in the archives, are based on our implementation of the data reduction methods as described in Howell et al. (2011) and Horch et al. (2012). Speckle reconstructions scale the output image maximum value to 1.0 at the center of the image. All of the images presented in this paper (except Nova V906 Car and Eros which are shown using just a red intensity color map) have their brightest pixel scaled to 1.0 and are presented on a min-max log scale using the MATLAB “jet” colormap (Figure 2).
Figure 2: The MATLAB “jet” color map used for the images presented in this paper.
@@
Fourier speckle deconvolution techniques are computationally efficient but require a PSF model power spectrum usually obtained via routine observation of stars taken from the HR or HD catalogs that are known or assumed to be single based on past spectroscopic or imaging observations. These PSF standards are assigned to each target by the observing team prior to the scheduled observations. They are observed near in time to the targets and with similar sky locations. Each PSF standard observation requires an additional $\sim 3$ minutes of observing time. Modern image reconstruction methods that are based on blind deconvolution techniques and can reach deeper contrasts plus provide more accurate astrophysical results are beginning to be implemented for speckle image reconstructions. (e.g., Howell et al., 2024).
@@
Figure 3 presents a typical Gemini speckle imaging result for a point source (the star TOI-5873), and the discovery of a very close stellar companion that is 1.6 magnitudes fainter than the primary. Note the 180-degree autocorrelation quadrant ambiguity that occurs when stars have comparable brightness (Horch et al., 2012). During data reduction, a “best determination”, using all sets of images taken of the target plus bispectrum analysis, is made to assign the correct location. The astrometric precision for binary stars observed at Gemini yields stellar separations to $\pm$1 mas and $\pm$1 degree in position angle (Lester et al., 2021). Photometric magnitude differences for stars in multiple systems typically have uncertainties of $\pm$0.25 magnitudes, increasing by 2$\sim$ 2$\times$ 2 × for very close or wide ($>$0.9”) pairs (e.g., Howell et al., 2011; Horch et al., 2012). Raw and fully reduced archival data from Zorro and ‘Alopeke can be found in the Gemini Observatory (Hirst and Cardenes, 2017) [^3] https://archive.gemini.edu/searchform/ZORRO,
@@
Figure 3: A typical speckle imaging result for an 8th magnitude star (TOI-5873). The red and blue curves show the $5\sigma$ contrasts achieved as a function of angular separation, from the diffraction limit out to 1.2 arcseconds. The star symbol shows the properties of the companion, and the inset shows the reconstructed image for the 832 nm observation; a scale bar is included for reference. The speckle observations revealed that the target star is binary, with a very close (0.14 arcsecond separation) companion that is 1.6 magnitudes fainter than the primary at position angle 176 degrees. See §3 for details.
@@ 6 A Panoply of Scientific Applications and Results
In this section, we discuss various scientific targets that have been observed using observations from ‘Alopeke and/or Zorro. Most of these observations were obtained as test cases or engineering studies (unless otherwise noted) and as such, many have not been published previously. These studies allowed researchers to assess the potential for using speckle imaging to accomplish their scientific goals, and have formed the basis for a number of stand-alone proposals targeting the detailed study of single objects, or larger sample collection to enable more global results.
@@ 6.1 Stars
Stars have been the predominant targets of the speckle imaging campaigns. In this section, we explore the varied results from these observations.
@@ 6.1.1 Exoplanet-Hosting Stars
Several large observing programs aimed at surveying exoplanet host stars have been in place at Gemini for many years. These programs have targeted transiting exoplanet candidates from Kepler, K2, TESS, and other NASA missions and have observed many thousands of stars (e.g., Lester et al., 2021; Howell et al., 2021a, b; Clark et al., 2022; Matson et al., 2025). Such surveys aim to identify close-in stellar companions that either induce false positive signals or contaminate the light curves of planets, leading to underestimated planetary radii (Ciardi et al., 2015; Furlan and Howell, 2017). Figure 3 is an example of an 8th magnitude TESS exoplanet candidate host star observed at Gemini South with Zorro. Observations of exoplanet host stars continue for missions such as TESS, Roman, Ariel, and the Habitable Worlds Observatory. As the Roman space telescope prepares for launch, microlensing observations will reach a new high as numerous free-floating planets will be discovered via observations pointed toward the Galactic bulge. Speckle observations will play an important role in the resolution of the target star lens, as was done for the Rsimilar-$\sim$18 Kojima-1Lb microlens event (Figure 4).
Figure 4: The microlens source Kojima-1Lb observed with ‘Alopeke at Gemini North. This SDSS $i$ band image was obtained using 35 minutes of on-source time and detected both the lens and source separated by 0.058 arcsec with a magnitude difference of 3.7. The very close (blended in the reconstructed image) companion shows a 180-degree ambiguity with the true source location being at P.A.=110 degrees. North is up and East to the left in the image.
@@ 6.1.2 Stellar Multiplicity
Surveys investigating the multiplicity of various types of stars have been carried out at the Gemini Observatory. These include low-mass stars (e.g., Winters et al., 2021; Clark et al., 2022), halo stars, and higher-order multiplicity in known and metal-poor binaries (e.g., Hartman et al., 2022; Mendez et al., 2025). Figure 5 shows a faint likely L-dwarf companion to an M star identified by Deacon and Hambly (2007), as well as a newly discovered, planet-hosting triple star system TOI-697 (Lester et al., 2021). High-resolution imaging of triple star systems has been proposed as a robust test of modified gravity (Manchanda et al., 2023).
Figure 5: Left: The late M star WISE J133734.15-371100.2 as observed by Zorro. This 9th magnitude star was observed for 30 minutes at 832 nm. Note the close companion at a separation of 0.6 arcsec that is 3 magnitudes fainter. At a distance of 34 pc, the two stars are 20 au apart and have masses of 0.23 and 0.09 M respectively. Right: The planet hosting star TOI-697 was discovered to be triple using Gemini speckle observations. The primary star harbors a 2.6 R planet in a 8.6 day orbit and is itself orbited by a close binary pair 5 magnitudes fainter at a separation of 1.15 arcsec (Armstrong et al., 2025). This 832 nm observation used 5 minutes of Gemini time.